OpenStudy (anonymous):

R1=1000.0Ω. R2=50.0Ω f=1560kHz

OpenStudy (anonymous):

Cmatch=444.7053 Lmatch=22.2352

OpenStudy (anonymous):

Cmatch = \[\frac{ \sqrt{R1*R2-R2^{2}} }{ 2*\pi*f*R1*R2 }\] Lmatch= \[\frac{ \sqrt{R1*R2-R2^{2}} }{ 2pif }\]